Factor the expression: x^4-20x^2+64 also, solveX^4-34X^2=-225
Answer by edjones(8007)

x^4-20x^2+64
=(x^2-16)(x^2-4)
=(x+4)(x-4)(x+2)(x-2)
.
x^4-34x^2=-225
x^4-34x^2+225=0
(x^2-25)(x^2-9)
(x+5)(x-5)(x+3)(x-3)
.
Ed